### v2

@@ Abstract
In this document, we discuss a multi-step approach to automated construction of a knowledge graph, for structuring and representing domain-specific knowledge from large document corpora. We apply our method to build the first knowledge graph of nuclear fusion energy, a highly specialized field characterized by vast scope and heterogeneity. This is an ideal benchmark to test the key features of our pipeline, including automatic named entity recognition and entity resolution. We show how pre-trained large language models can be used to address these challenges and we evaluate their performance against Zipf’s law, which characterizes human-generated natural language. Additionally, we develop a knowledge-graph retrieval-augmented generation system that combines large language models with a multi-prompt approach. This system provides contextually relevant answers to natural-language queries, including complex multi-hop questions that require reasoning across interconnected entities.
Index Terms:
Retrieval Augmented Generation, Knowledge Graph, Nuclear Fusion, Information Extraction, Large Language Models.
@@ I Introduction
As the area of research continues to expand, the need for an efficient system of data management and knowledge retrieval has become increasingly critical [3, 4, 5]. Addressing this challenge requires not only improving data accessibility but also enabling seamless interaction between human reasoning and machine processing. An interconnected system for online data, that can bridge this gap, holds significant appeal to many disciplines beyond nuclear fusion energy, as it has been clearly illustrated in [6, 7]. To address this need, new data models have been developed based on graph structures [8], and ontology frameworks [9].
In a graph, knowledge is stored in nodes (entities) linked together by edges. Nodes encapsulate concepts, objects, or nouns relevant to the domain, while edges define relationships between them. These relationships can carry semantic meaning like in subject-relationship-object triplets, or represent simpler connections such as the co-occurrence of two entities within the same text.
Ontology models, on the other hand, define a vocabulary of terms, their meanings, and organize them in hierarchical classes connected by well defined relationships. These two approaches represent complementary data models that, when used together, can enhance the accuracy and richness of the final Knowledge Base (KB), e.g., [10].
The development of a KB of fusion energy, which organizes information into structured categories and representations, will accelerate the development of more ”FAIR” data sharing, disseminating and referencing [3, 4]. Over time, such a system will foster collaborations between all partners in a fusion supply chain, helping to de-risk the design, development and construction of fusion power plants, while also reducing the barriers to understanding between field experts, funding agencies and policy makers.
In this work, we present an automated approach to generate a graph-based KB or Knowledge Graph (KG) of nuclear fusion energy starting from a large corpora of scientific documents and by leveraging the inference power of pre-trained Large Language Models (LLMs). Additionally, we developed a KG Retrieval-Augmented Generation (KG-RAG) machine. This combines the advanced linguistic capabilities of modern generative AI with domain-specific knowledge derived from previously unseen sources. The KG-RAG enables the effective retrieval of domain-specific information and the generation of accurate answers to user queries helping minimizing model hallucinations, i.e., the production of plausible but fictitious responses.
@@ II Related works
The creation of a KB begins with the process of Information Extraction (IE) from a text corpus. To automate the learning of linguistic features, different trainable algorithms have been developed over the years. These include approaches based on statistical analysis [11, 12], Bayesian inference [13, 14], and machine learning algorithms, e.g., [15, 16]. Early methods often relied on precompiled heuristic rules and human supervision to extract linguistic patterns e.g., [11] as well as on hand-annotated training sentences [17]. Other models employed ontologies to define classes and relationships of interest along with hand-labelled training texts, e.g., [15, 18]. For these kinds of Natural Language Processing (NLP) tasks, manual efforts remain a significant bottleneck, especially when dealing with large text corpora.
An essential task in IE is Named Entity Recognition (NER), which identifies entities specific to a particular domain within a given text. This process is often followed by Relation Extraction (RE), which establishes relationships between entities. Recurrent neural networks have been extensively used to accomplish these tasks, e.g., [19] and [20]. However, recent advancements in AI-powered NLP, particularly with LLMs, have enabled machines to assimilate vast amounts of text and process it with capabilities once believed to be uniquely human. Pre-trained LLMs have demonstrated good IE performance across disciplines, in both zero- and few-shot scenarios [21, 22, 23]. While their performance tends to decline when handling complex tasks [24] it can be recovered by breaking down high-level objectives into simpler, more manageable sub-tasks, each addressed through targeted prompts[25, 26].
The significant advancements, made over the years in the field of IE, have allowed the creation of several cross-disciplinary online KBs. Some of which were manually curated by their respective communities, as in the case of Freebase KG. Others, like DBpedia, were generated automatically from web sources such as Wikipedia [27]. To the best of our knowledge, a KB dedicated to nuclear fusion energy is not included in any of these databases, despite its societal relevance and the increasing interest shown by policymakers and the public. To address this gap, we embarked on the creation of the first KG for nuclear fusion energy.
@@ III-A Automated construction of a KG
A schematic of the pipeline for the automated construction of a KG is shown in figure 2. The workflow consists in the following steps: Data Acquisition (DAQ), NER, entity resolution, KG construction and RE.
Figure 2: Workflow for the automated creation of a KG. The first layer consists in the data acquisition. This involves scoping, to identify relevant data sources, formatting, to structure the data in JSON format. The second layer involves NER. This step is achieved through prompt-driven processing with a LLM. Following NER, a second LLM-assisted step performs entity resolution. This generates a set of rules for standardizing entities. The following step integrates the NER output and resolution rules to construct the KG. Entities that frequently appear together in the same context are paired. For each pair, the LLM extracts a relationship linking the two entities.
The DAQ layer involves two steps:
  • Scoping. This step identifies relevant data sources, which is accomplished by using an API to access the online archive of scholarly publications at www.lens.org.
  • Formatting. In this step, the retrieved data is structured in JSON format.
@@
We downloaded 8358 scientific abstracts from journal articles related to nuclear fusion energy and associated metadata such as: authors, keywords, publication year and scholarly citation count. The search process involved looking up specific patterns in article titles. A summary of the dataset is provided in Table I, along with the list of search patterns used. This approach may include a small number of articles that are not strictly related to nuclear fusion energy but match the search criteria1118 out of 335 in the Deuterium-Tritium sample. To filter out irrelevant abstracts, we implemented a data validation step using an LLM and a keyword-based filter. In this study, we have used Llama3.1 or LLama 3.3 by Meta, deployed on Amazon Web Services (AWS) equipped with 405 and 70 billion parameters, respectively 222meta.llama3-1-405b-instruct-v1:0 and meta.llama3-3-70b-instruct-v1:0.
Following the DAQ stage, the LLM was tasked with performing NER on the selected abstract text. To improve extraction accuracy, the text was split into individual sentences to reduce input length.
Previous works provided a definition of entity in their prompts effectively capturing the scope of their studies. Nevertheless, prompting for NER can be accomplished in few- or zero-shot mode depending on whether examples of entities are passed to the prompt or not, [21, 23, 25, 26]. In the present work, we do not provide explicit examples of entities but we do guide the extraction process by suggesting a list of category types to which entities belong. Although this method requires minimal user input, it allows to categorize entities in standard, cross-model communities (or groups) that may be used during the retrieval stage. To achieve optimal categorization, we have chosen category types that are as exclusive as possible, while also being broad enough to encompass a significant portion of relevant entities. Further refinement of these category types may be accomplished by running another NER pass with an updated LLM prompt. A list of the category types used in this work is given in the appendix. The accuracy of the classification process, that assigns each entity to a category type in the list, was estimated by comparing randomly selected entities and their LLM-assigned categories against the corresponding manually assigned ones. We acknowledge the limitations of this evaluation, as it depends on the specific choice of category types. Additionally, this estimate likely represents an upper bound on classification accuracy, given that some of the categories are not entirely exclusive and an entity may be correctly assigned to multiple categories. We found that the LLM correctly classified $\simeq$94of the entities  [^3] The accuracy was seen to decrease for LLMs with a smaller amount of parameters such as LLama3.1 8B.
Following the NER stage, the entity resolution step is performed which allows to resolve duplicates, acronyms and other inconsistencies among entities. During this process, entities are automatically standardized by using singular forms and lowercase spellings (when appropriate), while entities consisting solely of numbers are excluded. Additionally, this step removes entities inferred by the LLM that are not explicitly mentioned in the text. Part of entity resolution involves an additional LLM prompting pass. To this aim, a dictionary of top-ranked entities is created from the NER data and embedded into a prompt along with the simple instruction to identify acronyms with their expanded forms. Similar instructions are used to standardize chemical elements. The LLM then produces a list of substitutions required to transform entities into their standardized forms. This list is parsed during the graph creation stage ensuring consistency and reducing noise in the final knowledge graph.
@@
The entity resolution effectiveness can be appreciated by looking at the plots in figure 1. These show the frequency of the 500 top-ranked single word entities fitted by the Zipf law, before and after de-noising stage.
@@
The Zipf law is an empirical principle often observed in NLP and other areas of linguistics. It states that in a given corpus, the frequency $(f)$ of any word is inversely proportional to its rank $(r)$ in the frequency table: $f=C/r$. After entity resolution, the frequencies of some entities change resulting in a better fit to Zipf’s law.
@@
For the KG construction we have used the Neo4j framework. Each resolved entity is represented by an unique node in the graph. The graph structure is summarized in table II and includes nodes for abstracts, first authors, publication years, and key words, as well as individual sentences and entities. These nodes are interconnected through a set of predefined relationships, also summarized in table II and illustrated in figure 3:
@@
CC Entity/Entity weight
@@
Figure 3: Example of KG accordingly to the graph architecture used in this work. In this diagram, (Entity 1, RELATIONSHIP, Entity 2) form a triplet. Triplets have both [CC] and semantic relationships, see text for more details.
[…]
@@ III-B KG-RAG
Figure 4: Workflow of the KG-RAG. The user input is a question processed by a LLM that generates a Cypher query. A keyword search through the KG retrieves relevant documents (sentences) and triplets. These go through a sorter-filter module that yields a list of top-ranked sentences and triplets. Finally, the LLM generates an accurate answer to the user’s query.
The automated search of the KG for retrieval and generation purposes is accomplished by using the pipeline illustrated in figure 4. In the first step, the user submits a question and specifies the number of documents to retrieve. A first pass of LLM extracts relevant entities from the user’s question and embeds them into a Cypher query. Cypher is a declarative query language specifically designed for working with graph databases. Figure 5 shows two examples of LLM-generated Cypher queries generated from user inputs expressed in natural language. We considered both single-hop and multi-hop queries that can be answered by using information from a single node or multiple interconnected nodes, respectively.
We perform full-text search over the entire corpus to identify sentences that match entities extracted from the user’s query by the LLM. The retrieved sentences are then ranked in descending order of semantic similarity to the original question by a dedicated sorter module. Similarity scores are computed using the scalar product between the embedded vector representations of each sentence and the query. Sentence embeddings were generated using the ll-mpnet-base-v2 model. This maps a sentence into a 768 dimensional dense vector space and can be used for tasks like clustering or semantic search  [^5] https://huggingface.co/sentence-transformers/all-mpnet-base-v2. Finally, the top-k abstracts containing the most relevant sentences are selected and incorporated into the prompt provided to the LLM for answer generation.
The retriever also searches the graph for triplets, i.e., subject-relationship-object. These are then filtered based on two criteria: they must contain entities recognized by the LLM in the user’s query or they must include verbs with matching lemma to that in the query. If no match is found, the system sorts all the triplets by the total occurrence of the subject-object entities. Finally, triplets are passed to the LLM for answer generation.
@@ IV Results
In this study, we accomplished the automated construction of a KG of nuclear fusion energy that comprehensively represents the domain knowledge contained in 8358 scientific abstracts. To the best of our knowledge, this is the first KG in the domain of nuclear fusion energy. The graph structure is summarized in tables II and contains 108811 nodes and 718335 links.  [^6] The KG is available upon request: andrea.loreti@stfc.ac.uk
Assessing the performance of a RAG (or KG-RAG) system is complex, as it requires evaluating both retrieval accuracy and the faithfulness and correctness of generated responses, e.g., [28] and [29]. The interplay between RAG components, and the potential influence of the KG, make an exhaustive study of RAG performances beyond the scope of this work. However, we conducted two evaluation experiments as follows. First, we randomly selected $10$ abstracts and used ChatGPT to generate a specific question for each, simulating queries from field experts. We then examined the KG-RAG output to determine whether the corresponding abstracts were among the retrieved sources. In this experiment, the correct abstract was consistently ranked as the top result.
@@
Next, we generated a set of $50$ questions using a system prompt designed to simulate queries from different kinds of personas  [^7] We used Ragas: https://docs.ragas.io/en/v0.1.21/index.html. These questions were less specific, typically shorter, and contained typos or colloquial phrases. The retrieval accuracy for this broader set was $\simeq 50\%$. However, in all but one case, the retrieved documents remained highly relevant to the corresponding queries.
@@
For what is concerned with the generative process, this primarily consists in summarizing the retrieved top-ranked documents and does not add extra information from knowledge acquired during the pre-training, unless explicitly prompted to do so. If, on the one hand, these guardrails may limit the creative scope of AI tools, they also help to prevent LLMs from hallucinating and producing fictitious contextual answers, thus addressing a well-known issue affecting LLMs, e.g., [31]. While this answer generation can be very accurate, it lacks abstraction capabilities. For instance, the tasks of summarizing the content of the whole KG or extracting information within a specific time range require mutli-hop capabilities that retrieve information beyond the content of the documents themselves. KG-RAG offers a powerful way to achieve this level of abstraction by leveraging the entities and relationships stored in the graph. For example, by utilizing the predefined relationships outlined in table II and the LLM-generated semantic relationships, it has been possible to provide KG-RAG with an initial layer of abstraction, as it can be seen in examples 2 and 3 below.
@@
    User: ”How does the presence of high-density carbon impurities affect the ignition and burn of deuterium-tritium (DT) plasma, and what are the implications for inertial confinement fusion (ICF)?”
@@
    Sources:
    Abstract 1 [32]. One of the main concerns about the ignition and burn of deuterium-tritium DT plasma is the negative effect of impurities that can get into the thermonuclear fuel during target compression on the ignition of an inertial confinement fusion ICF target…
    Abstract 2 [33]. Accurate knowledge of the equation of state EOS of deuterium-tritium DT mixtures is critically important for inertial confinement fusion ICF. Although the study of EOS is an old topic, there is a longstanding lack of global accurate EOS data for DT within a unified theoretical framework… In the present work, we perform extensive state-of-the-art ab initio quantum Langevin molecular dynamics simulations to obtain EOS data for DT mixtures at densities from 0.1 gcm3 to 2000 gcm3 and temperatures from 500 K to 2000 eV, which are relevant to ICF processes…
    Abstract 3 [34]: Knowledge of the dynamics and concentration of impurities in the plasma core is crucial to the development of a deuterium-tritium DT tokamak fusion reactor…
@@
    User: ”Can you summarize the content of the database?”
    System: The data set appears to be a collection of sentences related to nuclear physics, specifically focusing on topics such as fusion reactions, plasma, and neutron interactions. The sentences describe various relationships between different entities, including deuterium, tritium, neutrons, and other particles, as well as concepts like ignition, implosion, and radiation.
    The data set seems to be a compilation of information from various sources, possibly research papers or scientific articles, and covers a range of topics, including:
    1. Fusion reactions and plasma physics
    2. Neutron interactions and scattering
    3. Ignition and implosion processes
    4. Radiation and heat transfer
    5. Particle physics and nuclear reactions
    The sentences are written in a formal, technical tone, suggesting that the data set is intended for an audience with a background in physics or a related field. Overall, the data set provides a comprehensive overview of various concepts and relationships in nuclear physics, but its purpose and intended use are unclear without further context.

### v3

@@ Abstract
In this document, we discuss a multi-step approach to automated construction of a knowledge graph, for structuring and representing domain-specific knowledge from large document corpora. We apply our method to build the first knowledge graph of nuclear fusion energy, a highly specialized field characterized by vast scope and heterogeneity. This is an ideal benchmark to test the key features of our pipeline, including automatic named entity recognition and entity resolution. We show how pre-trained large language models can be used to address these challenges and we evaluate their performance against Zipf’s law, which characterizes human natural language. Additionally, we develop a knowledge-graph retrieval-augmented generation system that uses multiple prompts with large language models to provide contextually relevant answers to natural-language queries, including complex multi-hop questions requiring reasoning across interconnected entities.
@@ I Introduction
As the area of research continues to expand, the need for an efficient system of data management and knowledge retrieval has become increasingly critical [29, 33, 25]. Addressing this challenge requires not only improving data accessibility but also enabling seamless interaction between human reasoning and machine processing.
Retrieval-Augmented Generation (RAG) has recently emerged as a promising method to address such challenges, e.g., [18]. Standard RAG systems employ embeddings to perform semantic similarity search and retrieving relevant documents from text corpora. While effective for many tasks, embeddings often capture similarity rather than logical relevance and struggle with multi-hop reasoning, e.g., [31]. These limitations can be mitigated by incorporating data models based on graph structures [3] or ontology frameworks [24]. These can enhance the reasoning capabilities of RAG architectures by providing more structured representations of the underlying Knowledge Base (KB).
The development of a KB of fusion energy, will accelerate the development of more ”FAIR” data sharing, disseminating and referencing [29, 33]. Over time, such a system will foster collaborations between all partners in a fusion supply chain, helping to de-risk the design, development and construction of fusion power plants, while also reducing the barriers to understanding between field experts, funding agencies and policy makers.
In this work, we present a simple yet effective automated approach to generate a graph-based KB or Knowledge Graph (KG) of nuclear fusion energy starting from a large corpora of scientific documents and by leveraging the inference power of pre-trained Large Language Models (LLMs). Additionally, we developed a KG Retrieval-Augmented Generation (KG-RAG) machine. This combines the advanced linguistic capabilities of modern generative AI with domain-specific knowledge derived from previously unseen sources. Our study shows the potential of KG-RAG systems to enhance information elicitation and retrieval in domains characterized by complex and interconnected data sources.
@@ II Related works
The creation of a KB begins with the process of Information Extraction (IE) from a text corpus. To automate the learning of linguistic features, different trainable algorithms have been developed over the years. These include approaches based on statistical analysis [27, 9], Bayesian inference [11, 7], and machine learning algorithms, e.g., [6, 4]. Early methods often relied on precompiled heuristic rules and human supervision to extract linguistic patterns e.g., [27] as well as on hand-annotated training sentences [10]. Other models employed ontologies to define classes and relationships of interest along with hand-labelled training texts, e.g., [6, 30]. For these kinds of natural language processing tasks, manual efforts remain a significant bottleneck, especially when dealing with large text corpora.
An essential task in IE is Named Entity Recognition (NER), which identifies entities specific to a particular domain within a given text. This process is often followed by Relation Extraction (RE), which establishes relationships between entities. Recurrent neural networks have been extensively used to accomplish these tasks, e.g., [35] and [20]. More recently, pre-trained LLMs have demonstrated good IE capabilities across disciplines, in both zero- and few-shot scenarios [1, 32, 12]. While their performance tends to decline when handling complex tasks [19] it can be recovered by breaking down high-level objectives into simpler, more manageable sub-tasks, each addressed through a targeted prompt [2, 5].
Over the years, various cross-disciplinary online KBs have been developed, e.g., Freebase or DBpedia [17]. To the best of our knowledge, none of these KBs focuses on nuclear fusion energy, despite its societal relevance and the increasing interest shown by policymakers and the public. To address this gap, we embarked on the creation of the first KG for nuclear fusion energy.
@@ III-A Automated construction of a KG
In a graph, knowledge is stored in nodes linked together by edges, see for instance figure 3. Nodes encapsulate entities relevant to the domain while edges define relationships between them. For the automatic construction of our KG we used the pipeline shown in figure 2. This is designed to be domain-agnostic and consists in the following: Data Acquisition (DAQ), NER, entity resolution, KG construction and RE.
Figure 2: Workflow for the automated construction of a KG, see text for more details.
DAQ
this stage comprises data scoping and formatting. First, relevant data sources are identified by accessing the archive of scholarly publications at www.lens.org via dedicated API. The retrieved data are then structured in JSON format.
@@
We downloaded 8358 scientific abstracts from journal articles related to nuclear fusion energy and associated metadata such as: authors, keywords, publication year and scholarly citation count. Abstracts were selected as they capture core terminology relevant for domain-level knowledge extraction. Articles were identified using title-based search patterns (Table I). Irrelevant abstracts  [^1] The searching criteria included a small number of non-relevant papers, 8 out of 335 in the Deuterium–Tritium sample. were filtered using an LLM-based validation step and a keyword filter.
Following the DAQ stage, the LLM performed NER on the selected texts.
NER
can be accomplished in few- or zero-shot mode depending on whether examples of entities are passed to the prompt or not, [1, 12, 2, 5]. In the present work, we proposed a set of category types that are as exclusive as possible while covering a large portion of relevant entities (see Appendix). While this introduces a degree of domain specificity it also enables standardized, cross-model categorization that may be used during the retrieval stage. Classification accuracy is estimated through human assessment of LLM-assigned categories on randomly selected entities. This estimate depends on the specific choice of category types. It also represents an upper bound on accuracy, since an entity may be validly assigned to multiple categories. We used Llama3.1 or Llama 3.3 by Meta, deployed on Amazon Web Services (AWS) equipped with 405 and 70 billion parameters, respectively 222meta.llama3-1-405b-instruct-v1:0 and meta.llama3-3-70b-instruct-v1:0. These models correctly classified $\lesssim$90% of the entities. The accuracy was seen to decrease to $\lesssim 70$% for smaller LLMs such as Llama3.1 8B. Notably, the landscape of LLMs has evolved rapidly since the beginning of this study, with the emergence of both compact dense models and sparse mixture‑of‑experts architectures that achieve comparable text‑processing performance to much larger dense models at significantly reduced computational costs, e.g., Phi‑3, Gemma 2, and DeepSeek‑V2/V3.
Entity resolution
the main sources of error in NER are entity hallucinations (LLM-inferred entities not present in the text), missed entities, and misclassifications. To reduce these errors, we narrowed the LLM’s focus by splitting text into individual sentences. We then applied entity resolution to standardize entity forms (converting to singular or lowercase where appropriate [^3] The spaCy software package was used for this task), exclude numeric-only entities, and remove hallucinations. As part of entity resolution, we performed an additional LLM pass to expand acronyms and standardize chemical element names (this step may require user supervision). Entity resolution improves consistency and reduces redundancy in the final KG.  [^4] For instance, it allowed to reduce the number of nodes from 4330 to 3632, and interlinks from 31076 to 25004 in a KG constructed from 334 abstracts.
@@
The effect of entity resolution can be appreciated by looking at the plots in figure 1. These show the frequency of the 500 top-ranked single word entities fitted by the Zipf’s law, before and after de-noising stage (entity resolution).
@@
The Zipf’s law is an empirical principle observed in linguistics. It states that in a given text corpus, the frequency $(f)$ of any word is inversely proportional to its rank $(r)$ in the frequency table: $f=C/r$. Applying Zipf’s law provides a quantitative benchmark to assess the naturalness and linguistic plausibility of the extracted entities. Deviations from its trend may reveal noise, or redundancy in the NER. After entity resolution, the frequency distribution of our extracted entities aligns better with Zipf’s law.
KG construction
@@
for this step we have used the Neo4j framework. The graph structure is summarized in table II and includes nodes for abstracts, first authors, publication years, and key words, as well as individual sentences and entities. These nodes are interconnected through a set of predefined relationships, also summarized in table II and illustrated in figure 3:
@@
$C$ Entity/Entity weight
@@
Figure 3: Example of KG accordingly to the graph architecture used in this work. In this diagram, (Entity 1, RELATIONSHIP, Entity 2) form a triplet. Triplets have both [$C$] and semantic relationships, see text for more details.
The link $\text{CC}]$ establishes the co-occurrence of two entities in the same sentence by assigning a weight to this relationship given by the number of times this co-occurrence happens.
Relation extraction
beyond the $\text{CC}]$ link, semantic relationships, such as predicative or causal connections, are inferred via a LLM pass. These are introduced only for highly recurrent nodes identified by filtering entities outside a three-standard-error range in bond strength and edge count. [^5] Sentences containing both candidate nodes are ranked by cosine similarity, and the top six are used by the LLM to infer relationships.
A qualitative inspection of the LLM‑inferred relationships reveals two main limitations. First, filtering entities by recurrence does not always guarantee semantically meaningful triplets, as highly frequent entities may correspond to generic terms that do not require explicit relational grounding (e.g., {optimization, applied_to, stellarator}). Second, extracted relations often take the form of full sentences rather than concise predicative or causal links, despite explicit prompting. While this behaviour does not imply factual inaccuracies, it introduces redundancy and reduces compactness. We suggest that these effects could be mitigated through stricter prompt constraints or post‑processing filters.
@@ III-B KG-RAG
Figure 4: Workflow of the KG-RAG, see text for more details.
The automated search of the KG for retrieval and generation purposes is accomplished by using the pipeline in figure 4. In the first step, the user submits a question and specifies the number k of top documents to retrieve. A first LLM pass extracts relevant entities and embeds them into a Cypher query. This is a language designed for working with graph databases. Figure 5 shows two LLM-generated Cypher queries generated from user inputs. We consider both single-hop and multi-hop queries that can be answered by using information from a single node or multiple interconnected nodes, respectively.
We perform full-text search to find sentences matching entities from the user’s query. These are ranked by semantic similarity using a sorter module, which computes cosine similarity between sentences and query embeddings  [^6] Embeddings are generated with the model https://huggingface.co/sentence-transformers/all-mpnet-base-v2, that maps text into a 768-dimensional space.. The top-k abstracts containing the most relevant sentences are then used to prompt the LLM for answer generation.
The retriever also searches the graph for triplets (subject–relationship–object). These are filtered to include only those containing entities from the user’s query or verbs with matching lemmas. If no match is found, triplets are ranked by subject–object frequency. The selected triplets are then passed to the LLM for answer generation.
@@ IV Results
In this study, we accomplished the automated construction of a KG of nuclear fusion energy that comprehensively represents the domain knowledge contained in 8358 scientific abstracts. To the best of our knowledge, this is the first KG in the domain of nuclear fusion energy. The graph structure is summarized in tables II and contains 108811 nodes and 718335 links (for the dataset of categorized entities see [23]).
Assessing the performance of a RAG (or KG-RAG) system requires evaluating both retrieval accuracy and the faithfulness and correctness of generated responses, e.g., [28] and [36]. An exhaustive study of RAG performances is beyond the scope of this work. However, we conducted two evaluation experiments. First, we randomly selected $10$ abstracts and used ChatGPT to generate one question each, simulating queries from field experts. We then checked whether the right abstract appeared among the retrieved sources. In this experiment, the correct abstract was consistently ranked as the top result.
@@
Next, we generated a set of $50$ questions using a system prompt designed to simulate queries from different kinds of personas  [^7] We used Ragas: https://docs.ragas.io/en/v0.1.21/index.html. These questions were less specific, shorter, and contained typos or colloquial phrases. The retrieval accuracy for this broader set was $\simeq 50\%$. However, in all but one case, the retrieved documents remained highly relevant to the corresponding queries showing the system capabilities to consistently surface relevant information.
@@
For what is concerned with the generative process, this primarily consists in summarizing the retrieved top-ranked documents and does not add extra information from knowledge acquired during the pre-training, unless explicitly prompted to do so, see example 1. If, on the one hand, these guardrails may limit the creative scope of generative AI, they also help to prevent LLMs from hallucinating and producing fictitious contextual answers, thus addressing a well-known issue affecting these models, e.g., [37]. Although accurate, answers generated by summarizing retrieved relevant documents may lack abstraction. Tasks such as summarizing the entire knowledge graph or extracting time-specific information often require multi-hop reasoning that goes beyond the content of individual retrieved documents. In a simple RAG system, achieving this level of abstraction can be difficult, as retrieval is limited to the top-k documents selected via embedding-based similarity. In a KG-RAG framework, this abstraction can be achieved by explicitly leveraging graph relationships. For instance, by using the links in (Table II) and LLM-generated semantic relations, KG-RAG gains an initial abstraction layer, as shown in examples 2 and 3.
@@
    User
    : ”How does the presence of high-density carbon impurities affect the ignition and burn of deuterium-tritium (DT) plasma, and what are the implications for inertial confinement fusion (ICF)?”
@@
    Sources: [26], [15] , [16].
@@
    User
    : ”Can you summarize the content of the database?”
    System: The data set appears to be a collection of sentences related to nuclear physics, specifically focusing on topics such as fusion reactions, plasma, and neutron interactions. The sentences describe various relationships between different entities, including deuterium, tritium, neutrons, and other particles, as well as concepts like ignition, implosion, and radiation. The data set seems to be a compilation of information from various sources, possibly research papers or scientific articles, and covers a range of topics, including…